### 01 Feb 2024 v1

@@ Appendix A Msss matyices
$\Delta_{2(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (62)
@@
×$\times{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
@@
$\Delta_{(2+i)(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}$ (63)
@@
In the same way, the CP-odd neutral scalars mass matrix is that:
@@
$M_{P}^{2}=\left(\begin{array}[]{ccccc}M_{\sigma_{d}\sigma_{d}}^{2}&M_{\sigma_{d}\sigma_{u}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{\sigma_{u}\sigma_{d}}^{2}&M_{\sigma_{u}\sigma_{u}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Im}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{1}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{2}^{c})^{\Im}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Im}(\tilde{\nu}_{3}^{c})^{\Im}}^{2}\\ \end{array}\right)$ (69)
$M_{\sigma_{d}\sigma_{d}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (70)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{66},$
$M_{\sigma_{d}\sigma_{u}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}$ (71)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{67},$
$M_{\sigma_{u}\sigma_{u}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (72)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{77},$
$M_{\sigma_{d}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}$ (73)
$+2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{i}^{c}}+\Delta_{6(7+i)},$
$M_{\sigma_{u}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}$ (74)
$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{i}^{c}}+\Delta_{7(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\Im}(\tilde{\nu}_{i}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{i}^{c}}}}-(A_{\kappa}\kappa)_{ijk}\Re(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{k}^{c}}$ (75)
$+4\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}+\Delta_{(7+i)(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\Im}(\tilde{\nu}_{j}^{c})^{\Im}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{u}^{2}+v_{d}^{2})+\Delta_{(7+i)(7+i)}.$ (76)
@@
$\Delta_{66}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (77)
@@
$\Delta_{67}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (78)
×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{77}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (79)
@@
$\Delta_{6(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (80)
×$\times{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{7(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (81)
×$\times{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{(7+i)(7+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (82)
@@
$\Delta_{66}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}})^{2}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (83)

### 06 Mar 2024 v2

@@ Acknowledgements.
$\Delta_{2(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (62)
@@
×$\times{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
@@
$\Delta_{(2+i)(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u})^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}$ (63)
@@
In the same way, the $CP$-odd neutral scalars mass matrix is that
@@
$M_{P}^{2}=\left(\begin{array}[]{ccccc}M_{\sigma_{d}\sigma_{d}}^{2}&M_{\sigma_{d}\sigma_{u}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{\sigma_{d}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{\sigma_{u}\sigma_{d}}^{2}&M_{\sigma_{u}\sigma_{u}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{\sigma_{u}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Im}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Im}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}\sigma_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}\sigma_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}(\tilde{\nu}_{1}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}(\tilde{\nu}_{2}^{c})^{\text{Im}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Im}}(\tilde{\nu}_{3}^{c})^{\text{Im}}}^{2}\\ \end{array}\right)$ (69)
$M_{\sigma_{d}\sigma_{d}}^{2}\!\!\!\!\!$ $=\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (70)
$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{66},$
$M_{\sigma_{d}\sigma_{u}}^{2}\!\!\!\!\!$ $=-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}$ (71)
$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{67},$
$M_{\sigma_{u}\sigma_{u}}^{2}\!\!\!\!\!$ $=\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (72)
$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{77},$
$M_{\sigma_{d}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}$ (73)
$+2\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{i}^{c}}+\Delta_{6(7+i)},$
$M_{\sigma_{u}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}$ (74)
$-2\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{i}^{c}}+\Delta_{7(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\text{Im}}(\tilde{\nu}_{i}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{i}^{c}}}}-(A_{\kappa}\kappa)_{ijk}\text{Re}(e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})\upsilon_{\nu_{k}^{c}}$ (75)
$+4\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}v_{u}+\Delta_{(7+i)(7+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\text{Im}}(\tilde{\nu}_{j}^{c})^{\text{Im}}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{u}^{2}+v_{d}^{2})+\Delta_{(7+i)(7+i)},$ (76)
@@
$\Delta_{66}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (77)
@@
$\Delta_{67}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (78)
×$\times{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{77}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta)^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$ (79)
@@
$\Delta_{6(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (80)
×$\times{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{7(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (81)
×$\times{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{(7+i)(7+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{(-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}})^{2}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (82)
@@
$\Delta_{66}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{(-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}})^{2}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$ (83)